OpenStudy (anonymous):

Write an equation for the line in point/slope form and slope/intercept form that has the given condition. 2. Passing through (-2, 4) and (-1, -1).

OpenStudy (anonymous):

@iGreen @SarahMotionless

OpenStudy (anonymous):

@confluxepic @Ilovecake

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

Im gonna be honest I don't know how to do it.

OpenStudy (anonymous):

hold up let me try to find a lesson

OpenStudy (anonymous):

Okay cool

OpenStudy (anonymous):

Heres an example Point slope form is: y-y1 = m(x-x1)...the subscripts are to denote the use of the points So you're formula would be: y - 0 = -2(x+2) the positive because two negatives become a positive slope-intercept form is: y=mx+b where m is the slope and b is the y intercept. The problem is the y intercept is not given to you directly so you must solve for it using the slope and points: y = mx+b 0 = -2(-2) + b 0 = 4 + b -4 = b Now use the y-intercept to set up the equation: y = -2x-4

OpenStudy (anonymous):

Okay ummm im not sure what to do

OpenStudy (anonymous):

@amirreza1870 can you help me?

OpenStudy (anonymous):

If you have two points you can find the slope of their line with this formula \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] then use one of your point in this formula\[y-y _{1}=m(x-x _{1})\]